OpenStudy (anonymous):

What is the LCD of 3k/40 and k/18 ?

OpenStudy (anonymous):

least common multiple of 40 and 18 is 360

OpenStudy (anonymous):

\(40=2\times 20,18=2\times 9, 2\times 20\times 9=360\)